An aluminum piston 4 in. in diameter that moves in the cylinder of a cast iron engine block must have a clearance diameter of 0.003 in. when hot. The temperature rises from 70 degrees to 160 degrees Farenheit. What should the clearance between the bore diameter and the piston be if the linear coefficient of expansion per unit length (inch) per degree Farenheit is 0.0000124 for aluminum and 0.00000655 for cast iron when the temperature is 70 degrees F?
Thank you, in advance for your time and attention to this matter.
Mary Miller

Answer by Alan3354(69443)

An aluminum piston 4 in. in diameter that moves in the cylinder of a cast iron engine block must have a clearance diameter of 0.003 in. when hot. The temperature rises from 70 degrees to 160 degrees Farenheit. What should the clearance between the bore diameter and the piston be if the linear coefficient of expansion per unit length (inch) per degree Farenheit is 0.0000124 for aluminum and 0.00000655 for cast iron when the temperature is 70 degrees F?
----------------
Al: alpha = 1.24E-5
FE: alpha = 6.55E-6
------
If the Al piston is 4 in diameter at 70F:
Its diameter at 160F = 4 + (160-70)*4*1.24E-5
= 4 + 446.4E-5 = 4 + 4.465E-3
= 4.004465 inches
--------------
If the bore is 4.003 at 70F, it changes to:
4.003 + 4.003*90*6.55E-6
= 4.003 + 2359.8E-6 = 4.003 + 2.36E-3
= 4.00536 inches
----------------
4.00536 - 4.004465 = 0.000895", less than 1/1000 of an inch.
------------------
In reality, the expansion of the bore hole in the solid block is more complex. The iron expands to make the bore larger, but the surrounding structure also contributes to "squeezing" the hole, so it's not a linear function.